# Moment-based parameter inference with error guarantees for stochastic reaction networks

### v1

@@ Abstract
Inferring parameters of models of biochemical kinetics from single-cell data remains challenging because of the uncertainty arising from the intractability of the likelihood function of stochastic reaction networks. Such uncertainty falls beyond current error quantification measures, which focus on the effects of finite sample size and identifiability but lack theoretical guarantees when likelihood approximations are needed. Here, we propose an inference method for stochastic reaction networks with nonlinear and rational propensities at steady state that provides bounds on the parameters via convex optimisation over sets constrained by moment equations and moment matrices. Our approach takes observations from the stochastic reaction network and forms moment intervals, which are then used to constrain parameters through convex sets. The bounds on the parameters contain the true parameters under the condition that the moment intervals contain the true stationary moments, thus providing uncertainty quantification and error guarantees. Our approach does not need to predict moments and distributions for given parameters (i.e., it avoids solving or simulating the forward problem), and hence circumvents intractable likelihood computations or computationally expensive simulations. We demonstrate its use for uncertainty quantification, data integration and prediction of latent species statistics through synthetic data from common nonlinear biochemical models including the Schlögl model, the toggle switch and post-transcriptional regulation.
@@ I Introduction
Biological processes in single cells are influenced by fluctuations in the timing of biochemical reactions McAdams and Arkin (1997); Raser and O’shea (2005); Taniguchi et al. (2010). Single-cell analysis has become an essential tool to quantify variability across living cells and holds the promise to unravel the mechanisms underlying cellular functions through advances in flow cytometry, microscopy, and omics approaches. Yet the quantitative understanding of these experiments necessitates mechanistic models that account for the observed cell-to-cell variabilityShahrezaei and Swain (2008); Thomas, Popović, and Grima (2014); Gorin, Vastola, and Pachter (2023). The chemical master equation (CME) provides such a mathematical description governing the probability distributions of stochastic reaction networksGillespie (1977); Goutsias and Jenkinson (2013); Anderson and Kurtz (2015); Schnoerr, Sanguinetti, and Grima (2017); Wilkinson (2018); Kuntz et al. (2021). Making accurate predictions using these models requires parameter inference from experimental data and quantifying the uncertainty in these predictions, a challenging task prompting more efficient and precise inference methodsSchnoerr, Sanguinetti, and Grima (2017); Loskot, Atitey, and Mihaylova (2019); Warne, Baker, and Simpson (2019).
@@
In recent years, a few authors have obtained theoretically guaranteed bounds on the stationary moments of stochastic reaction networks (Kuntz et al., 2016; Sakurai and Hori, 2017; Ghusinga et al., 2017; Dowdy and Barton, 2018a, b; Kuntz et al., 2019; hori2020modeling). These approaches rely on convex optimisation of sets constrained by moment equations, a set of equations involving the stationary moments and the reaction rate parameters, and positive semidefinite constraints on the moments. By optimising over these convex sets with a given set of rate parameters, one can obtain upper and lower bounds on the stationary moments. These bounds provide error guarantees for predicting moments, unlike approximation methods based on system size expansion or moment closure. Whether similar approaches could be utilised to provide error bounds for parameter inference has remained unexplored heretofore.
In this work, we present an approach to inference in stochastic reaction networks with unobserved species, which provides upper and lower bounds on parameters in stationary conditions. Our approach takes intervals of the stationary moments as inputs and constrains the parameter space by considering the moment equations and moment matrix constraints to formulate an optimisation problem. The optimisation is turned into a convex semidefinite program (SDP) through a hierarchy of outer approximations with a unique global optimum that allows efficient computation using available SDP solvers. If the input moment intervals contain the true moments, our approach guarantees to provide bounds on the parameters that include the true values. Our method extends to bound moments of unobserved species. A visual summary of our approach is shown in Fig. 1.
@@
Figure 1: Flowchart of inference using moment constraints. Given a stochastic reaction network model (potentially with unobserved species), we use a series of constraints involving stationary moments and reaction rate parameters to form a constrained set on them. The mathematical expressions of these constraints can be found in (12) and (III.1). From count data, we obtain bootstrap intervals on the moments and input these into the set. By introducing additional variables to replace the non-linear terms in the set, we have a semi-definite program (SDP) over which we optimise with respect to a parameter (or an unobserved moment) to obtain upper and lower bounds.
II Notation and background: The CME and the Stationary Moment Equations
@@
The CME consists of a set of differential equations of the probability distribution $P_{t}\coloneqq(P_{t}(\bm{x}))_{\bm{x}\in\mathcal{X}}$ over time $t$ and is commonly formulated in the form
$\frac{dP_{t}(\bm{x})}{dt}=P_{t}Q(\bm{x})\coloneqq\sum_{\bm{y}\in\mathcal{X}}P_{t}(\bm{y})q(\bm{y},\bm{x}),$ (2)
$q(\bm{x},\bm{y})\coloneqq\sum_{r=1}^{R}a_{r}(\bm{x})(\mathbb{I}_{\bm{x}+\bm{v}_{r}}(\bm{y})-\mathbb{I}_{\bm{x}}(\bm{y})),$ (3)
@@
where $b_{r}(\bm{x})$ are multivariate polynomials in $\bm{x}$. For rational propensities(Kuntz et al., 2019), we can always rewrite them in the form
$a_{r}(\bm{x})=k_{r}\frac{b_{r}(\bm{x})}{h(\bm{x})},\quad r=1,\ldots,R.$ (6)
Since the polynomial propensities are a special case of rational ones with $h\equiv 1$, we will only discuss the rational ones from now on. We denote the highest degree of $b_{r}$ by $d_{b}$ and the degree of $h$ by $d_{h}$. The common denominator in all reactions $h$ can be written as
@@
for a multi-index ${\bm{\alpha}}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{S})$.
Before introducing the moment equations, we first define the set of stationary distributions of the CME as any probability $\pi\coloneqq(\pi(\bm{x}))_{\bm{x}\in\mathcal{X}}$ distribution that solves
$\pi Q(\bm{x})=0,\quad\forall\bm{x}\in\mathcal{X}.$ (10)
For a stationary distribution $\pi$, we denote the expectation of a function $f\colon\mathcal{X}\rightarrow\mathbb{R}$ as $\langle f\rangle_{\pi}\coloneqq\sum_{\bm{x}\in\mathcal{X}}f(\bm{x})\pi(\bm{x})$ such that the corresponding vector of rational moments is denoted as
$\bm{y}^{\bm{\alpha}}\coloneqq\left(\left\langle\frac{\bm{x}^{\bm{\alpha}}}{h(\bm{x})}\right\rangle_{\pi}\right)^{\top}_{{\bm{\alpha}}\in\mathbb{N}^{S}}.$ (11)
and we say this rational moment has order $|\bm{\alpha}|$. We define the moment vector $\bm{y}_{:d}$ as the set of rational moments with order less than or equal to $d$. For a given order $d$ and number of species $S$, the length of the moment vector $\bm{y}_{:d}$ is $N_{d}\coloneqq\binom{S+d}{S}$. These moment vectors satisfy the stationary moment equations:
$\left\langle Q\bm{x}^{\bm{\alpha}}\right\rangle_{\pi}=\bm{k}^{\top}A^{d}_{\bm{\alpha}}\bm{y}_{:d}=0,$ (12)
@@
$A^{d}_{\bm{\alpha}}(r,m)\coloneqq\sum_{\begin{subarray}{c}\bm{\gamma}:\ \bm{l}_{m}-d_{b}\leq|\bm{\gamma}|\leq|\bm{\alpha}|-1\end{subarray}}\binom{\bm{\alpha}}{\bm{\gamma}}b_{r,\bm{l}_{m}-\bm{\gamma}}\bm{v}_{r}^{\bm{\alpha}-\bm{\gamma}},$ (13)
if $|\bm{l}_{m}|\leq|\bm{\alpha}|+d_{b}-1$ and zero otherwise, which depends on the numerator $b_{r}(\bm{x})$ of the propensity functions and the stoichiometric matrix $\bm{v}_{r}$ but not on the reaction parameter $k_{r}$ (see Appendix A for a derivation). The constant $|\bm{\alpha}|+d_{b}-1$ is the order of the highest order moment with non-zero coefficient involved in the moment equation, which we refer to as the order of the moment equation.
@@ III Theoretical results: Bounds on Rate Parameters via Optimisation
In this section, we introduce our approach to defining sets of parameters constrained by the moment equations and moment matrices, and we use outer approximations of these sets to form convex optimisation problems. We consider two cases of data availability: the complete data case (Sec. III.1), in which joint-moments of all species $X_{1},\ldots,X_{S}$ are measured in the experiment, and the partial data case (Sec. III.2), in which moments of only a subset of species are observed.
III.1 Complete Data Case
If the true moments are known, one can solve a sufficient number of moment equations in (12) directly for the rate parameters (assuming that the parameters are uniquely identifiable). In practice, the true moments are unknown, and we have to use estimates instead. If we were to substitute the true moments with the estimators $\hat{\bm{y}}$, the moment equations could easily be violated due to sampling errors. We therefore consider moment intervals $(\hat{\bm{y}}_{l},\hat{\bm{y}}_{u})$ that are assumed to contain the true moments and reflect their statistical uncertainty. Such intervals can be obtained by bootstrapping the data, for example, but our approach is not limited to this method.
@@ III.1 Complete Data Case
The first inequality is the moment interval data containing the true moment. The second row is the moment equations satisfied by the true moments and parameters for any $d\geq d_{b}-1$. The third constraint comes from the fact that $\sum_{\bm{x}\in\mathcal{X}}\pi(\bm{x})=1$ and thus the rational moments satisfy the condition
@@
assuming $N_{d}\geq d_{h}$ and $\bm{h}_{:N_{d}}\coloneqq(\bm{h}^{\top},0,\ldots,0)^{\top}$ is a zero-padded version of $\bm{h}\coloneqq(h_{\bm{\alpha}})_{|\bm{\alpha}|\leq d_{h}}$, the coefficients in $h(\bm{x})$. In the case of polynomial moments ($h\equiv 1$), this condition becomes $y_{0}=1$. The final constraint arises from the fact that $\bm{y}$ are moments of a measure over the positive quadrants, with positive semidefinite inequalities associated with the moment matrices(Bertsimas and Sethuraman, 2000; Lasserre, 2008) defined as
@@
$Z\coloneqq\bm{y}\bm{k}^{\top}=(\bm{z}_{1}\ \cdots\ \bm{z}_{R})\in\mathbb{R}_{+}^{(N_{d}\times R)},$ (18)
@@
$\tilde{\xi}^{d}=\left\{{\begin{array}[]{l}\bm{k}\in\mathcal{K}\\ \bm{y}\in{\mathbb{R}_{+}^{N_{d}}}\\ Z\in\\ \ \ \mathbb{R}_{+}^{(N_{d}\times R)}\end{array}}{\begin{array}[]{|ll}k_{j}\hat{\bm{y}}_{l}\leq\bm{z}_{j}\leq k_{j}\hat{\bm{y}}_{u},&j=1,\ldots,R\\ \text{trace}(A^{d}_{\bm{\alpha}}Z)=0,&|{\bm{\alpha}}|\leq d-d_{b}+1\\ \bm{z}_{j}^{\top}\bm{h}_{:N_{d}}=k_{j},&j=1,\ldots,R\\ M_{d}^{s}(\bm{y})\succeq 0,&s=0,\ldots,S\\ ZC^{\top}=\bm{y}\bm{c}^{\top}\\ \text{rank}(Z)=1\end{array}}\right\}$ (20)
where the condition $ZC^{\top}=\bm{y}\bm{c}^{\top}$ follows from $C\bm{k}=c$.
@@
$\zeta^{d}=\left\{{\begin{array}[]{l}\bm{k}\in\mathcal{K}\\ \bm{y}\in{\mathbb{R}_{+}^{N_{d}}}\\ Z\in\\ \ \ \mathbb{R}_{+}^{(N_{d}\times R)}\end{array}}{\begin{array}[]{|ll}k_{j}\hat{\bm{y}}_{l}\leq\bm{z}_{j}\leq k_{j}\hat{\bm{y}}_{u},&j=1,\ldots,R\\ \text{trace}(A^{d}_{\bm{\alpha}}Z)=0,&|{\bm{\alpha}}|\leq d-d_{b}+1\\ \bm{z}_{j}^{\top}\bm{h}_{:N_{d}}=k_{j},&j=1,\ldots,R\\ M_{d}^{s}(\bm{y})\succeq 0,&s=0,\ldots,S\\ ZC^{\top}=\bm{y}\bm{c}^{\top}\\ \end{array}}\right\}.$ (21)
@@
III.2 Partial Data Case
A straightforward generalisation relevant in practice is to assume that not all species are observed, so that only partial information and intervals for moments associated with the observed species are available. We split the moment vector $\bm{y}$ into the observed moments ${\bm{y}}^{obs}$ and the unobserved moments $\bm{y}^{unobs}$, which involves the unobserved species

### v2

@@ Abstract
Inferring parameters of models of biochemical kinetics from single-cell data remains challenging because of the uncertainty arising from the intractability of the likelihood function of stochastic reaction networks. Such uncertainty falls beyond current error quantification measures, which focus on the effects of finite sample size and identifiability but lack theoretical guarantees when likelihood approximations are needed. Here, we propose a method for the inference of parameters of stochastic reaction networks that works for both steady-state and time-resolved data and is applicable to networks with non-linear and rational propensities. Our approach provides bounds on the parameters via convex optimisation over sets constrained by moment equations and moment matrices by taking observations to form moment intervals, which are then used to constrain parameters through convex sets. The bounds on the parameters contain the true parameters under the condition that the moment intervals contain the true moments, thus providing uncertainty quantification and error guarantees. Our approach does not need to predict moments and distributions for given parameters (i.e., it avoids solving or simulating the forward problem), and hence circumvents intractable likelihood computations or computationally expensive simulations. We demonstrate its use for uncertainty quantification, data integration and prediction of latent species statistics through synthetic data from common non-linear biochemical models including the Schlögl model and the toggle switch, a model of post-transcriptional regulation at steady state, and a birth-death model with time-dependent data.
@@ I Introduction
Biological processes in single cells are influenced by fluctuations in the timing of biochemical reactions McAdams and Arkin (1997); Raser and O’shea (2005); Taniguchi et al. (2010). Single-cell analysis has become an essential tool to quantify variability across living cells and holds the promise to unravel the mechanisms underlying cellular functions through advances in flow cytometry, microscopy, and omics approaches. Yet the quantitative understanding of these experiments necessitates mechanistic models that account for the observed cell-to-cell variabilityShahrezaei and Swain (2008); Thomas, Popović, and Grima (2014); Dattani and Barahona (2017); Gorin, Vastola, and Pachter (2023). The chemical master equation (CME) provides such a mathematical description governing the probability distributions of stochastic reaction networksGillespie (1977); Goutsias and Jenkinson (2013); Anderson and Kurtz (2015); Schnoerr, Sanguinetti, and Grima (2017); Wilkinson (2018); Kuntz et al. (2021). Making accurate predictions using these models requires parameter inference from experimental data and quantifying the uncertainty in these predictions, a challenging task prompting more efficient and precise inference methodsSchnoerr, Sanguinetti, and Grima (2017); Loskot, Atitey, and Mihaylova (2019); Warne, Baker, and Simpson (2019).
@@
In recent years, a few authors have obtained theoretically guaranteed bounds on the stationary moments (Kuntz et al., 2016; Sakurai and Hori, 2017; Ghusinga et al., 2017; Dowdy and Barton, 2018a, b; Sakurai and Hori, 2018a; Kuntz et al., 2019; Hori, 2020; Sakurai and Hori, 2022) and the transient moments (Dowdy and Barton, 2018b; Sakurai and Hori, 2018b; Holtorf and Barton, 2023) of stochastic reaction networks. These approaches rely on convex optimisation of sets constrained by moment equations, a set of equations involving the moments and the reaction rate parameters, and positive semidefinite constraints on the moments. By optimising over these convex sets with a given set of rate parameters, one can obtain upper and lower bounds on the moments. These bounds provide error guarantees for predicting moments, unlike approximation methods based on system size expansion or moment closure. Whether similar approaches could be utilised to provide error bounds for parameter inference has remained unexplored heretofore.
In this work, we present an approach to inference in stochastic reaction networks with unobserved species, which provides upper and lower bounds on parameters. Our approach can be applied to data observed at steady state as well as time trajectory data (i.e. transient data). It takes intervals of the moments as inputs and constrains the parameter space by considering the moment equations and moment matrix constraints to formulate an optimisation problem. The optimisation is turned into a convex semidefinite program (SDP) through a hierarchy of outer approximations with a unique global optimum that allows efficient computation using available SDP solvers. If the input moment intervals contain the true moments, our approach guarantees to provide bounds on the parameters that include the true values. Our method extends to bound moments of unobserved species. A visual summary of our approach is shown in Fig. 1.
@@
Figure 1: Flowchart of inference using moment constraints. Given a stochastic reaction network model (potentially with unobserved species), we use a series of constraints involving the moments and the reaction rate parameters to form a constrained set on them. The mathematical expressions of these constraints can be found in Sec. II.3. From count data at steady state or time-resolved data, we obtain bootstrap intervals on the moments and input these into the set. By introducing additional variables to replace the non-linear terms in the set, we have a semi-definite program (SDP) over which we optimise with respect to a parameter (or an unobserved moment) to obtain upper and lower bounds.
II Notation and background: The CME and the Moment Equations
@@
II.1 Chemical Master Equation
The CME consists of a set of differential equations of the probability distribution $\bm{P}_{t}\coloneqq(P_{t}(\bm{x}))_{\bm{x}\in\mathcal{X}}$ over time $t$ and is commonly formulated in the form
$\frac{\mathrm{d}P_{t}(\bm{x})}{\mathrm{d}t}=\bm{P}_{t}Q(\bm{x})\coloneqq\sum_{\bm{y}\in\mathcal{X}}P_{t}(\bm{y})q(\bm{y},\bm{x}),$ (2)
$q(\bm{y},\bm{x})\coloneqq\sum_{r=1}^{R}a_{r}(\bm{y})(\mathbb{I}_{\bm{y}+\bm{v}_{r}}(\bm{x})-\mathbb{I}_{\bm{y}}(\bm{x})),$ (3)
@@ II.1 Chemical Master Equation
where $k_{r}$ is the rate constant and $b_{r}(\bm{x})$ are multivariate polynomials in $\bm{x}$. For rational propensities(Kuntz et al., 2019), we can always rewrite them in the form
$a_{r}(\bm{x})=k_{r}\frac{b_{r}(\bm{x})}{h(\bm{x})},\quad r=1,\ldots,R,$ (6)
where $h$ is the common denominator in all reactions. More specifically, we denote the highest degree of $b_{r}$ by $d_{b}$ and the degree of $h$ by $d_{h}$. Then $h$ can be written as
@@
for a multi-index ${\bm{\alpha}}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{S})$. We denote the coefficients in $h(\bm{x})$ by $\bm{h}\coloneqq(h_{\bm{\alpha}})_{|\bm{\alpha}|\leq d_{h}}$. Since the polynomial propensities are a special case of rational ones with $h\equiv 1$, we will only discuss the rational ones from now on.
II.2 Moment Equations
For a probability distribution $P_{t}$, we denote the expectation of a function $f\colon\mathcal{X}\rightarrow\mathbb{R}$ as $\langle f\rangle_{P_{t}}\coloneqq\sum_{\bm{x}\in\mathcal{X}}f(\bm{x})P_{t}(\bm{x})$. A raw moment is then given by
$\mu^{\bm{\alpha}}(t)\coloneqq\left\langle\bm{x}^{\bm{\alpha}}\right\rangle_{P_{t}}$ (10)
and the corresponding rational moment
$y^{\bm{\alpha}}(t)\coloneqq\left\langle\frac{\bm{x}^{\bm{\alpha}}}{h(\bm{x})}\right\rangle_{P_{t}}$ (11)
for a multi-index $\bm{\alpha}$, and we say each moment has order $|\bm{\alpha}|$. Let us define the moment vector $\bm{y}_{:d}(t)$ as the set of rational moments with order less than or equal to $d$. For a given order $d$ and number of species $S$, the length of the moment vector $\bm{y}_{:d}(t)$ is $N_{d}\coloneqq\binom{S+d}{S}$. The raw and rational moments are related by the linear identity
$\mu^{\bm{\alpha}}(t)={\bm{\bar{h}}_{\bm{\alpha}}}^{\top}\bm{y}_{:d}(t),$ (12)
for $d\geq d_{h}+|\bm{\alpha}|$, indicating that they provide the equivalent information, and $\bm{\bar{h}}_{\bm{\alpha}}\in\mathbb{R}^{N_{d}}$ is defined as
$\bm{\bar{h}}_{\bm{\alpha}}(\bm{x})\coloneqq\left\{\begin{array}[]{ll}h_{\bm{x}-\bm{\alpha}},&\text{if }\bm{x}\geq\bm{\alpha}\text{ and }|\bm{x}-\bm{\alpha}|\leq d_{h},\\ 0,&\text{otherwise.}\end{array}\right.$ (13)
Denoting the vector of rate constants as
$\bm{k}=(k_{1},k_{2},\ldots,k_{R})^{\top},$ (14)
the moment vectors satisfy the transient moment equations:
$\frac{\mathrm{d}\mu^{\bm{\alpha}}}{\mathrm{d}t}=\left\langle Q\bm{x}^{\bm{\alpha}}\right\rangle_{P_{t}}=\bm{k}^{\top}A^{d}_{\bm{\alpha}}\,\bm{y}_{:d}(t),$ (15)
@@ II.2 Moment Equations
$A^{d}_{\bm{\alpha}}(r,\bm{l})\coloneqq\sum_{\begin{subarray}{c}\bm{\gamma}:\ \bm{l}-d_{b}\leq|\bm{\gamma}|\leq|\bm{\alpha}|-1\end{subarray}}\binom{\bm{\alpha}}{\bm{\gamma}}b_{r,\bm{l}-\bm{\gamma}}\bm{v}_{r}^{\bm{\alpha}-\bm{\gamma}},$ (16)
if $|\bm{l}|\leq|\bm{\alpha}|+d_{b}-1$ and zero otherwise, which depends on the numerator $b_{r}(\bm{x})$ of the propensity functions and the stoichiometric matrix $\bm{v}_{r}$ but not on the reaction parameter $k_{r}$ (see Appendix A for a derivation). The constant $|\bm{\alpha}|+d_{b}-1$ is the order of the highest order moment with a non-zero coefficient involved in the moment equation, which we refer to as the order of the moment equation.
II.3 Stationary and Generalised Transient Moment Equations
Here, we seek to construct algebraic constraints between the moments of the reaction network and its parameters. The simplest case is obtained in steady state where the vector of rational moments $\bm{y}_{:d}=\lim_{t\to\infty}\bm{y}_{:d}(t)$ satisfies the stationary moment equations:
$\bm{k}^{\top}A^{d}_{\bm{\alpha}}\,\bm{y}_{:d}=0,$ (17)
for any moment vector with $d\geq|\bm{\alpha}|+d_{b}-1$, which are obtained by setting the time-derivative in (15) to zero. Different algebraic relations hold in the time-dependent case. To this end, we define the generalised rational moments as a summary statistic obtained over the time interval $[0,T)$:
[…]
where $\rho\in\mathbb{R}$ is a constant. One can similarly define $\bm{\bar{y}}_{:d}(\rho)$ corresponding to $\bm{y}_{:d}(t)$ to denote the set of generalised moments up to order $d$. The generalised transient moment equation satisfied by these moments are
$\mu^{\bm{\alpha}}(T)-e^{\rho T}\mu^{\bm{\alpha}}(0)+\rho\bar{\mu}^{\bm{\alpha}}(\rho)=\bm{k}^{\top}A^{d}_{\bm{\alpha}}\,\bm{\bar{y}}_{:d}(\rho)$ (19)
for $|\bm{\alpha}|\leq d-d_{b}+1$, which are obtained by multiplying (15) with $e^{\rho(T-t)}$ and using partial integration(Dowdy and Barton, 2018b). Denoting $\bm{y}_{:d}(t)$ as $\bm{y}(t)$ for $t=0,T$ and $\bm{\bar{y}}_{:d}(\rho)$ as $\bm{\bar{y}}(\rho)$, and using (12), the generalised transient moment equation can then be written as
${\bm{\bar{h}}_{\bm{\alpha}}}^{\top}(\bm{y}(T)-e^{\rho T}\bm{y}(0)+\rho\bm{\bar{y}}(\rho))=\bm{k}^{\top}A^{d}_{\bm{\alpha}}\,\bm{\bar{y}}(\rho)$ (20)
for $|\bm{\alpha}|\leq d-d_{h}-d_{b}+1$. In the following, we will demonstrate how these relations can be used to infer parameters compatible with different moment data.
@@ III Theoretical results: Bounds on Rate Parameters via Optimisation
In this section, we introduce our approach to defining sets of parameters constrained by the moment equations and moment matrices, and we use outer approximations of these sets to form convex optimisation problems. For steady-state data, we consider two cases of data availability: the complete data case (Sec. III.1), in which joint-moments of all species $X_{1},\ldots,X_{S}$ are measured in the experiment, and the partial data case (Sec. III.2), in which moments of only a subset of species are observed. Finally, in Sec. III.3, we extend our approach so that it can be applied when transient data are observed.
III.1 Complete Data Case at Steady State
If the true moments are known, one can solve a sufficient number of moment equations directly in (15) for the rate parameters (assuming that the parameters are uniquely identifiable). In practice, the true moments are unknown, and we have to use estimates instead. If we were to substitute the true moments with the estimators $\hat{\bm{y}}$, the moment equations could easily be violated due to sampling errors. We therefore consider moment intervals $(\hat{\bm{y}}_{l},\hat{\bm{y}}_{u})$ that are assumed to contain the true moments and reflect their statistical uncertainty. Such intervals can be obtained by bootstrapping the data, for example, but our approach is not limited to this method (see discussions in Sec V).
@@ III.1 Complete Data Case at Steady State
The first inequality is the moment interval data containing the true moment. Note that although we consider bounds on rational moments here, bounds on raw moments can equivalently be used via (12). The second row is the moment equations satisfied by the true moments and parameters for any $d\geq d_{b}-1$. The third constraint comes from the fact that $\sum_{\bm{x}\in\mathcal{X}}\pi(\bm{x})=1$ and thus the rational moments satisfy the condition
@@
assuming $N_{d}\geq d_{h}$ and $\bm{h}_{:N_{d}}\coloneqq(\bm{h}^{\top},0,\ldots,0)^{\top}$ is a zero-padded version of $\bm{h}$, the coefficients in $h(\bm{x})$. In the case of polynomial moments ($h\equiv 1$), this condition becomes $y_{0}=1$. The final constraint arises from the fact that $\bm{y}$ are moments of a measure over the positive quadrants, with positive semidefinite inequalities associated with the moment matrices(Bertsimas and Sethuraman, 2000; Lasserre, 2008) defined as
@@
$Z\coloneqq\bm{y}\bm{k}^{\top}=(\bm{z}_{1}\ \cdots\ \bm{z}_{R})\in\mathbb{R}_{+}^{N_{d}\times R},$ (25)
@@
$\tilde{\xi}^{d}=\left\{{\begin{array}[]{l}\bm{k}\in\mathcal{K}\\ \bm{y}\in{\mathbb{R}_{+}^{N_{d}}}\\ Z\in\\ \ \ \mathbb{R}_{+}^{N_{d}\times R}\end{array}}{\begin{array}[]{|ll}k_{j}\hat{\bm{y}}_{l}\leq\bm{z}_{j}\leq k_{j}\hat{\bm{y}}_{u},&j=1,\ldots,R\\ \text{trace}(A^{d}_{\bm{\alpha}}Z)=0,&|{\bm{\alpha}}|\leq d-d_{b}+1\\ \bm{z}_{j}^{\top}\bm{h}_{:N_{d}}=k_{j},&j=1,\ldots,R\\ M_{d}^{s}(\bm{y})\succeq 0,&s=0,\ldots,S\\ ZC^{\top}=\bm{y}\bm{c}^{\top}\\ \text{rank}(Z)=1\end{array}}\right\}$ (27)
where the condition $ZC^{\top}=\bm{y}\bm{c}^{\top}$ follows from $C\bm{k}=\bm{c}$.
@@
$\zeta^{d}=\left\{{\begin{array}[]{l}\bm{k}\in\mathcal{K}\\ \bm{y}\in{\mathbb{R}_{+}^{N_{d}}}\\ Z\in\\ \ \ \mathbb{R}_{+}^{N_{d}\times R}\end{array}}{\begin{array}[]{|ll}k_{j}\hat{\bm{y}}_{l}\leq\bm{z}_{j}\leq k_{j}\hat{\bm{y}}_{u},&j=1,\ldots,R\\ \text{trace}(A^{d}_{\bm{\alpha}}Z)=0,&|{\bm{\alpha}}|\leq d-d_{b}+1\\ \bm{z}_{j}^{\top}\bm{h}_{:N_{d}}=k_{j},&j=1,\ldots,R\\ M_{d}^{s}(\bm{y})\succeq 0,&s=0,\ldots,S\\ ZC^{\top}=\bm{y}\bm{c}^{\top}\\ \end{array}}\right\}.$ (28)
@@
III.2 Partial Data Case at Steady State
A straightforward generalisation relevant in practice is to assume that not all species are observed so that only partial information and intervals for moments associated with the observed species are available. We split the moment vector $\bm{y}$ into the observed moments ${\bm{y}}^{obs}$ and the unobserved moments $\bm{y}^{unobs}$, which involves the unobserved species